Alex2007:

If f(x)=−2x−3 and g(x)=x2+5x, find the value. f(0)−7=

carlosbigbain:

To find the value of f(0) - 7, we can start by finding the value of f(0) and then subtracting 7 from it. Let's break it down step by step: The function f(x) is given as f(x) = -2x - 3. To find f(0), we substitute 0 for x in the function: f(0) = -2 * 0 - 3. Simplifying further: f(0) = 0 - 3. And the value of f(0) is: f(0) = -3. Now, we subtract 7 from f(0): f(0) - 7 = -3 - 7. Simplifying further: f(0) - 7 = -10. So, the value of f(0) - 7 is -10